OpenStudy (anonymous):

This is a probability question that I positive I got the answer but I want to make sure

OpenStudy (anonymous):

The thing is that it is really close to one another so I have no idea if it will be answer choice one or answer choice 3

OpenStudy (anonymous):

\[\frac{475}{500}\] vs \[\frac{475}{850}\] they don't look that close to me

OpenStudy (anonymous):

oops typo there, i meant \(\frac{475}{550}\)

OpenStudy (anonymous):

the prob of being male is 550/100

OpenStudy (anonymous):

the two event are not independent

OpenStudy (anonymous):

btw the 100 should a 1000